5) Find the point (L,y) on the graph of y= Vr nearest the point (1,0).
Answer #1

y = 5x suppose that (x,y) is nearest to the point (1,0). Distance between the two pintor V2-x)? H0-y)? Substitute y = 5x Dist. f(x) is minimum when x= 1/2 When f(x) is minimum, the distanie is minimum . Minimum distance: √(1-1½)² + 1 = V / 4 + 1/2 √3